Consider the following discrete probability distribution 15 22 P(X x) 0.14 0.38 0.27 0.21 xc 34 40 a. Is this a valid probability distribution? O Yes, because the probabilities add up to 1. O No, because the gaps between x values vary. b. What is the probability that the random variable X is less than 36? (Round your answer to 2 decimal places.) Probability c. What is the probability that the random variable X is between 12 and 30? (Round your answer to 2 decimal places.) Probability
d. What is the probability that the random variable X is greater than 17? (Round your answer to 2 decimal places.) Probability
Answer #1

b) P( X < 36) = P( X = 15) + P( X = 22) + P( X = 34) = 0.14 + 0.38 + 0.27 = 0.79

c)  P( 12 < X < 30 ) = P( X = 15) + P( X = 22) = 0.14 + 0.38 = 0.52

d) P( X > 17 ) = P( X = 22) + P( X = 34) + P( X = 40) = 0.38 + 0.27 + 0.21 = 0.86